As part of the new year celebrations, the town council decorated the street with banners hung across the tops of buildings. A section of the main street has two buildings 24 meters apart. The heights of the buildings are 13 meters and 20 meters. If the banners are hung tautly, what is the length, in meters, of the shortest banner joining the tops of these Iwo buildings? Write your answer in the space below.
Answer by mananth(16946)

BRM is a right angled triangle
BR^2= BM^2+MR^2
BR^2 = 24^2+7^2 =625
BR = sqet(625)= 25
25 m
.